Suppose that about 77% of graduating students attend their graduation. A group of 33 students is randomly chosen, and let X be the number of students who attended their graduation.

  1. What is the distribution of X? X ~ 
  2. ? B U N 
  3.  (,) 
  4. Please show the following answers to 4 decimal places.
  5. What is the probability that exactly 25 number of students who attended their graduation in this study? 
  6. What is the probability that less than 25 number of students who attended their graduation in this study? 
  7. What is the probability that at most 25 number of students who attended their graduation in this study? 
  8. What is the probability that between 20 and 26 (including 20 and 26) number of students who attended their graduation in this study? 
1
Expert's answer
2022-03-21T02:05:37-0400

Solution;

1.The distribution of X is

"x\approx B(n,p)=(33,0.77)"

2.

"P(x=25)=(\displaystyle_x^n)p^x(1-p)^{n-x}"

"P(x=25)=(\displaystyle _{25}^{33})(0.77)^{25}(1-0.77)^{33-25}=0.2579"

6.

"P(x<25)=\displaystyle{\sum_{x=0}^{24}}(\displaystyle_x^n)p^x(1-p)^{n-x}"

"P(x<25)=\displaystyle{\sum_{x=0}^{24}}(\displaystyle_{24}^{33})(0.77)^{24}(1-0.77)^{33-24}=0.3427"

7.

"P(x\leq25)=\displaystyle{\sum_{x=0}^{25}}(\displaystyle_x^n)p^x(1-p)^{n-x}"

"P(x\leq25)=\displaystyle{\sum_{x=0}^{24}}(\displaystyle_{25}^{33})(0.77)^{25}(1-0.77)^{33-25}=0.4997"

8.

"P(20\leq x\leq 26)=\displaystyle{\sum_{x=20}^{26}}(\displaystyle_x^n)p^x(1-p)^{n-x}"

"P(20\leq x\leq 26)=\displaystyle{\sum_{x=20}^{26}}(\displaystyle_x^{33})(0.77)^x(1-0.77)^{n-x}=0.6522"